OpenStudy (anonymous):

Two cars leave towns 400 kilometers apart at the same time and travel toward each other. One car's rate is 12 kilometers per hour more than the other's. If they meet in 2 hours, what is the rate of the faster car?

OpenStudy (anonymous):

@kevindragonlord @sangya21 @UltimateWitherLord @uri @Kit_Kat_Nat_<3 @Lori_The_Time_Lady_ @nelsonjedi can u help

OpenStudy (anonymous):

hmmm let me think

OpenStudy (anonymous):

I need to know at least the speed of one car.

OpenStudy (anonymous):

it says one car's rate is 12 kilometers @kevindragonlord

OpenStudy (anonymous):

@freckles @camper4834 can u helpo

OpenStudy (anonymous):

one of them are going 200 kph and the other is going 188kph

OpenStudy (anonymous):

actually wait one of them would be going 200kph and the other 212kph

OpenStudy (anonymous):

Deandre drove to the mountains last weekend. There was heavy traffic on the way there, and the trip took 8 hours. When Deandre drove home, there was no traffic and the trip only took 5 hours. If his average rate was 21 miles per hour faster on the trip home, how far away does Deandre live from the mountains? Do not do any rounding.

OpenStudy (anonymous):

400=2*(x+(x+12)) 400-24= 4x x = 94 x+12=106 kmph for the faster car check 400 = 2*94 + 2*106 400 = 188 + 212 400 = 400

OpenStudy (anonymous):

i got the answer already, can u help with this one @mathmath333 @mathstudent55 @uri @sangya21 Deandre drove to the mountains last weekend. There was heavy traffic on the way there, and the trip took 8 hours. When Deandre drove home, there was no traffic and the trip only took 5 hours. If his average rate was 21 miles per hour faster on the trip home, how far away does Deandre live from the mountains? Do not do any rounding.

OpenStudy (anonymous):

let x=driving speed to Deandre x+26=driving speed on return trip distance=speed*travel time .. 8x=5(x+21) 8x=5x+105 3x=105 x=35 x+21=46 driving speed to deandre=35 mph driving speed on return trip=46 mph

OpenStudy (anonymous):

your speed and time is given 8 hours and 35mph distance = 8 *35 = 280 miles

OpenStudy (anonymous):

A car passes a landmark on a highway traveling at a constant rate of 55 kilometers per hour. One hour later, a second car passes the same landmark traveling in the same direction at 63 kilometers per hour. How much time after the second car passes the landmark will it overtake the first car? Do not do any rounding.

OpenStudy (anonymous):

@sangya21 @mathmath333 @mathstudent55 @lizzyrockefeller can u help ^^